OpenStudy (anonymous):

given g(x)= 2x^2+2p+1, find g(p+2). simplify answer

OpenStudy (anonymous):

g(x)= 2x^2+2p+1, g(p+2)=2(p+2)^2+2p+1=2(p^2+4p+4)+2p+1=2p^2+10p+9